Dickinson Company has $12,020,000 million in assets. Currently half of these assets are financed with long-term debt at 10.1 percent and half with common stock having a par value of $8. Ms. Smith, Vice President of Finance, wishes to analyze two refinancing plans, one with more debt (D) and one with more equity (E). The company earns a return on assets before interest and taxes of 10.1 percent. The tax rate is 40 percent. Tax loss carryover provisions apply, so negative tax amounts are permissable.

Under Plan D, a $3,005,000 million long-term bond would be sold at an interest rate of 12.1 percent and 375,625 shares of stock would be purchased in the market at $8 per share and retired.

Under Plan E, 375,625 shares of stock would be sold at $8 per share and the $3,005,000 in proceeds would be used to reduce long-term debt.

b-3. Compute the earnings per share if return on assets increased to 15.1 percent. Round your answers to 2 decimal places.) Current Plan Plan E Earnings per s hare b-4. Which plan would be most favorable if return on assets increased to 15.1 percent? Consider the current plan and the two new plans. O Plan E O Plan D Current Plan c-1. If the market price for common stock rose to $10 before the restructuring, compute the earnings per share. Continue to assume that $3,005,000 million in debt will be used to retire stock in Plan D and $3,005,000 million of new equity will be sold to retire debt in Plan E. Also assume that return on assets is 10.1 percent. (Round your answers to 2 decimal places.) Current Plan Plan D Plan E Earnings per s hare
c-2. If the market price for common stock rose to $10 before the restructuring, which plan would then be most attractive? Plan D Current Plan Plan E
Answer #1

b - 3

Please see the table below. Please see the second column titled "Linkage". This column explain how each and every value is calculated. Figures in parenthesis are negative values. Shares repurchased and debt repaid appear as negative values. The final values appear in the last row. They had been highlighted in yellow colored cells. Those are your final answers for part b - 3.

Parameter

Linkage

Current Plan

Plan D

Plan E

Total Assets

A

12,020,000

12,020,000

12,020,000

Existing debt

B = 50% x A

6,010,000

6,010,000

6,010,000

Interest rate on existing debt

C

10.10%

10.10%

10.10%

Existing equity

D = 40% x A

6,010,000

6,010,000

6,010,000

Par value of shares

E

8

8

8

Existing numbers of shares outstanding

F = D / E

751,250

751,250

751,250

Restructuring

Fresh debt

G

-

3,005,000

(3,005,000)

Interest rate on fresh debt

H

12.10%

10.10%

Number of shares issued / (repurchased)

I

-

(375,625)

375,625

After restructuring

Interest expenses

J = B x C + G x H

607,010

970,615

303,505

Number of shares outstanding

K = F + I

751,250

375,625

1,126,875

EPS Calculation

ROA

L

15.10%

15.10%

15.10%

EBIT

M = L x A

1,815,020

1,815,020

1,815,020

[-] Interest expenses

N = J above

607,010

970,615

303,505

EBT

O

1,208,010

844,405

1,511,515

[-] Taxes

P = 40% x O

483,204

337,762

604,606

Net income

Q

724,806

506,643

906,909

EPS

Q / K

0.96

1.35

0.80

b - 4

Plan D has the highest EPS. Hence the correct answer is second option: Plan D

C - 1

We will have to rework the whole thing with ROA of 10.1% and equity issuance price of $ 10 / share

Nos. of new shares that will be issued or repurchased = 3,005,000 / 10 = 300,5000

Please see the table below. Last row in the table shows all your answers.

Parameter

Linkage

Current Plan

Plan D

Plan E

Total Assets

A

12,020,000

12,020,000

12,020,000

Existing debt

B = 50% x A

6,010,000

6,010,000

6,010,000

Interest rate on existing debt

C

10.10%

10.10%

10.10%

Existing equity

D = 40% x A

6,010,000

6,010,000

6,010,000

Par value of shares

E

8

8

8

Existing numbers of shares outstanding

F = D / E

751,250

751,250

751,250

Restructuring

Fresh debt

G

-

3,005,000

(3,005,000)

Interest rate on fresh debt

H

12.10%

10.10%

Number of shares issued / (repurchased)

I

-

(300,500)

300,500

After restructuring

Interest expenses

J = B x C + G x H

607,010

970,615

303,505

Number of shares outstanding

K = F + I

751,250

450,750

1,051,750

EPS Calculation

ROA

L

10.10%

10.10%

10.10%

EBIT

M = L x A

1,214,020

1,214,020

1,214,020

[-] Interest expenses

N = J above

607,010

970,615

303,505

EBT

O

607,010

243,405

910,515

[-] Taxes

P = 40% x O

242,804

97,362

364,206

Net income

Q

364,206

146,043

546,309

EPS

Q / K

0.48

0.32

0.52

c - 2

Since plan E has the highest EPS, Hence correct answer is last option showing Plan E.